If a man is capable of lifting a 131 lb stone on land, how heavy a stone could he lift under water? The specific gravity of the stone is 2.70.
Answer #1

We know that

Density = Mass/Volume

so, Mass = Density x Volume
Let Ds be the stone's density and Dw be the density of water
Ds/Dw = 2.7

Ds = 2.7*Dw
Let the Volume of the submerged stone be V
So,

131 = [DsV - Dw*V]*g
131/g = V[2.7Dw-Dw]

131/9.81= V*1.7Dw
V = 131/[9.81*1.7*Dw]
The weight of the submerged stone is then
DsV*g = 2.7*Dw*g*131/[g*1.7*Dw]

= 2.7*131/1.7

= 208.058 lbs